Find the length of the major and minor axes of an ellipse with the equation 16x2 + 25y2 + 32x - 150y = 159.
Answer by fractalier(6550)

16x^2 + 25y^2 + 32x - 150y = 159
rearrange terms and get
16x^2 + 32x + 25y^2 - 150y = 159
now factor out and complete the squares
16(x^2 + 2x + 1) + 25(y^2 - 6y + 9) = 159 + 16 + 225
16(x + 1)^2 + 25(y - 3)^2 = 400
(x + 1)^2 / 25 + (y - 3)^2 / 16 = 1
The major axis is 2a, here since a^2 = 25, it is 10...
The minor axis is 2b, here since b^2 = 16, it is 8...